# An accuracy-runtime trade-off comparison of scalable Gaussian process approximations for spatial data

### v1

@@ 2.4 Gaussian process approximations and choice of tuning parameters
We consider the following methods: Vecchia approximations, covariance tapering, modified predictive process (MPP) aka Fully Independent Training Conditional (FITC) approximations, full-scale approximations (FSA) combining predictive processes with tapering, multi-resolution approximations (MRA), the stochastic partial differential equations (SPDE)-based approximation of Lindgren et al. (2011b), fixed rank kriging (FRK), and the periodic embedding approach of Guinness (2019).
@@
The runtime and accuracy of the different approximation methods are determined by one or several tuning parameters such as the number of neighbors for Vecchia approximations and the number of knots for modified predictive process approximations. For every approximation, we use up to five different choices of tuning parameters. The selection of tuning parameter is done such that the different methods have similar run-times. The specific choices of the tuning parameters thus vary across datasets and are reported in Chapter 3. In the simulated experiments, the choice of the tuning parameters is done based on the time for evaluating the log-likelihood. In real-world applications, we instead use the total time for estimation and prediction on the (“interpolation”) test set. We also set an upper bound on the total runtime for every iteration in the experiments, including likelihood evaluation, estimation, and prediction, and we do not consider tuning parameter choices that exceed this limit. For simulated datasets with sample sizes $N=10^{\prime}000$ and $N=100^{\prime}000$, this limit is $10$ and $30$ minutes per repetition, respectively. For the real-world data, we use the following limits: $30$ minutes for the house price dataset, two hours for the MODIS 2016 dataset, three hours for the Laegern dataset, and four hours for the MODIS 2023 dataset.
@@
Vecchia approximations are applied to the observable process $y(\cdot)$ and not the latent process $f(\cdot)$, random orderings are used, nearest neighbors are determined using the Euclidean distance, and for prediction, the observed locations appear first in the ordering and we condition on the observed locations only. For FITC and FSA approximations, the knots are determined using the kmeans++ algorithm. For covariance tapering, the time for log-likelihood evaluation does not include the time required to perform a symbolic Cholesky decomposition for sparsity pattern since this needs to be done only once.
@@ 2.5 Software and hardware resources
We use the following software implementations for conducting our experiments. For Vecchia approximations, covariance tapering, FITC approximations, and full-scale approximations, we use the GPBoost library version 1.5.1 (https://github.com/fabsig/GPBoost, Sigrist et al. (2021)). All these methods are implemented in C++ in a comparable manner which makes their comparison as fair as possible. For the other four techniques, we use the following R packages: GPvecchia version 0.1.5 (https://github.com/katzfuss-group/GPvecchia) for the multi-resolution approximation, INLA version 24.6.27 (www.r-inla.org, Rue et al. (2009), Lindgren et al. (2011a)) and rSPDE version 2.3.3 (Bolin and Simas, 2023) for the SPDE approach, FRK version 2.3.0 for fixed rank kriging (Zammit-Mangion and Cressie, 2021), and npspec version 0.1.0 for periodic embedding (https://github.com/joeguinness/npspec). Parameter estimation is done with the default optimizers in the corresponding software packages.
@@ 3.1 Simulated data with $N=10^{\prime}000$
We first discuss the results for the simulated data with a sample of size $N=10^{\prime}000$ and an practical range of $0.2$. We use figures to present the results in the main article, the results in tabular form can be found in the appendix. The tuning parameters used for every method are reported in Table 1.
@@
In Figure 6, we present the absolute mean difference between the exact and approximate log-likelihood versus the evaluation time. The left panel shows the results when the evaluation is done at the true data-generating parameters, while the right panel shows the results using “wrong” parameters obtained by doubling the true parameters. Vecchia approximations clearly achieve the highest accuracy in computing the log-likelihood in both parameter settings. The difference to the exact log-likelihood decreases quickly with respect to time, approaching values close to zero even for small numbers of neighbors. The second most accurate approximation for a given computational budget is the FITC approximation followed by the MRA and FSA. Tapering is the method with the overall lowest accuracy, achieving a high accuracy only for relatively large taper ranges. The results for larger and smaller practical ranges of $0.5$ and $0.05$ are reported in Appendix C and Appendix D, respectively. Overall, we observe similar findings. Vecchia approximations are clearly the most accurate approximations. As expected, FITC approximations are more accurate for large range parameters compared to small ones, and vice versa for covariance tapering. Note that the SPDE-based approach is not included in the log-likelihood comparison as, to the best of our knowledge, the corresponding software implementation does not allow for evaluating the log-likelihood at user-defined covariance parameters. Further, FRK and and periodic embedding do not try to approximate a GP with a Matern covariance function, and it is thus pointless to analyze the accuracy of log-likelihood evaluations.
@@
Figure 6: Absolute mean difference between approximate and exact log-likelihood on simulated datasets with an practical range of $0.2$ of and a sample size $N=10^{\prime}000$. The true data-generating parameters are used on the left and “wrong” parameters obtained by doubling the true ones are used on the right.
@@
We next present the results concerning parameter estimation accuracy. In Figure 7, we show the bias and the MSE of estimates for the range, marginal variance, and error variance parameters, respectively, versus the time required to estimate these parameters. We again find that Vecchia approximations are clearly the most accurate with the lowest bias and MSE for most computational times and very small biases and MSEs even for small estimation times. The range and the variance parameters are overestimated by most methods. For tapering and FITC, some results with extreme values are excluded from the plots for better visibility of the other results. The full results in tabular form can be found in Appendix B. Note that we were not able to extract estimates for range and marginal variance parameters using the corresponding software implementation for the SPDE approach.
@@
Figure 7: Bias and MSE of the estimates for the range, marginal variance, and error term variance on simulated datasets with an practical range of $0.2$ and a sample size $N=10^{\prime}000$. The dashed lines correspond to the results from the exact calculations.
@@
In Figure 8, we report the prediction accuracy results including RMSEs, log-scores, and KL divergences to exact calculations for predicting the latent process at training locations, the “interpolation” test data, and the “extrapolation” test locations. We observe that Vecchia approximations are clearly the most accurate compared to all other methods regarding all evaluation metrics for most computational times. Overall, the MRA is the second most accurate method in this experiment, followed by the FITC and SPDE approximations. The FITC, FSA, MRA, and SPDE methods eventually achieve similar prediction accuracies as Vecchia approximations but for much larger computational times. We were not able to attain this level of accuracy with tapering in the computational times considered. Finally, FRK results in considerably less accurate predictions. For the “extrapolation” test set, we observe that the variability is much higher for all methods due to the uncertainty involved in making predictions outside of the training domain. Note that for FRK, the metrics calculated on the “extrapolation” set are not included in the figure as they were very large, i.e., very inaccurate. All results in tabular form can be found in Appendix B. In Appendix B, we additionally report the RMSE comparing the approximate predictive means and variance to the exact values.
@@
Figure 8: Average RMSE, log-score, and KL divergence between exact and approximate predictions on simulated datasets with an practical range of $0.2$ and a sample size $N=10^{\prime}000$. We distinguish between predicting the latent process at (i) training locations (left), (ii) interpolation test locations (center), and (iii) extrapolation locations (right). The dashed lines correspond to the results from the exact calculations.
@@
Results for experiments conducted on simulated datasets with practical ranges of 0.5 and 0.05 are reported in Appendices C and D, respectively. In both scenarios, Vecchia approximations consistently rank among the most accurate, particularly for short runtimes. The MRA shows similar behavior across the different ranges, though it tends to be slightly less accurate than Vecchia approximations. All techniques generally are more accurate for a larger range parameter. Some techniques, however, are more affected by the range parameter than others. A notable example is covariance tapering, which, as expected, is more accurate for small range parameters and less accurate for large ones. For the FITC approximation, the situation is reversed: it is more accurate for large range parameters and less accurate for small ones.
3.2 Simulated data with $N=100^{\prime}000$
Next, we present the results for the simulated data with a sample of size $N=100^{\prime}000$ and an practical range of $0.2$. The results for the smaller and larger ranges as well as all tabulated results can be found in the appendix. The tuning parameters used for every method are shown in Table 2.
Figure 9 displays the average approximate negative log-likelihood versus the evaluation time. As before, we consider the evaluation at the true data-generating parameters and using “wrong” parameters obtained by doubling the true parameters. Recall that, in this case, we do not consider the difference to the exact log-likelihood, as the latter could not be calculated in a reasonable amount of time. The results are very similar to the ones of the smaller sample size. The Vecchia approximation is again the most accurate method, with the maximal accuracy being achieved for relatively small runtimes, i.e., number of neighbors.
[…]
Concerning parameter estimation accuracy, Figure 10 reports the bias and MSE of estimates for the range, marginal variance, and error variance parameters, respectively, versus the time required to estimate these parameters. We again find very similar results as for the smaller sample size of $N=10^{\prime}000$. Vecchia approximations are clearly the most accurate with the lowest bias and MSE for most computational times and very small biases and MSEs even for small number of neighbors. Note that for tapering, some results with extreme values are again excluded from the plots for better visibility of the other results. The full results in tabular form can be found in Appendix E.
Figure 10: Bias and MSE of the estimates for the range, marginal variance, end error term variance on simulated datasets with an practical range of $0.2$ and a sample size N=100’000,
For assessing prediction accuracy, Figure 11 shows the RMSE and log-scores for predicting the latent process at training locations, the “interpolation” test data, and the “extrapolation” test data. We find that Vecchia approximations are the most accurate compared to all other methods regarding all evaluation metrics for most computational times. The SPDE-based approach results in the second most accurate predictions overall followed by the MRA and the FITC approximation. FRK and tapering generally generate the least accurate predictions. Note that the FITC approximation performs relatively well on the extrapolation test data. For FRK, the results have been excluded again from the extrapolation plots due to extreme and very inaccurate values; they can be found in Appendix 49 in tabular form.
Figure 11: Average RMSE and log-score on simulated datasets with an practical range of $0.2$ and a sample size N=100’000.
@@ 3.2 Simulated data with $N=100^{\prime}000$
The results for the experiments using simulated data with $N=100^{\prime}000$ and practical ranges of $0.5$ and $0.05$ are reported in Appendix F and Appendix G, respectively. Overall, the results are qualitatively very similar compared to the above-reported results.
3.3 House price dataset
In this subsection, we present the results for the house price dataset. Note that we do not use the periodic embedding approach for this dataset, as the locations are not arranged on a lattice. The tuning parameters used on this data set are reported in Table 3.
In Figure 12, we present the prediction accuracy results on the “interpolation” and “extrapolation” test sets of the housing dataset. Specifically, we plot the RMSE, the log-score and the CRPS versus the prediction time. Note that, since estimating the covariance parameters is a necessary first step on real-world data, we include the estimation time as part of the total time required to make predictions. In addition, we also report the results when doing predictions with exact calculations. To further compare the differences between exact and approximate calculations, Figure 13 shows the KL divergence between the exact and approximate predictive distributions versus the runtime on the two test sets.
Vecchia approximations are again the most accurate approximations having almost equal prediction accuracy metrics as exact calculations on both the interpolation and the extrapolation test sets even for small runtimes, i.e., small numbers of neighbors. This is confirmed by the very small KL divergences in Figure 13 of predictive distributions of Vecchia approximations compared to exact calculations. For the interpolation test set, full-scale approximations are second most accurate. For long runtimes, the prediction quality metrics are very similar across all approximations except for FRK. Note that for the extrapolation test data, the FITC and full-scale approximations with a small number of inducing points lead to more accurate predictions than an exact Gaussian process models. This means that simpler approximations can perform better than exact GPs for extrapolation. As above, the poor FRK scores on the “extrapolation” set had to be removed from the plots, but can be found in Appendix H.
In Figure 14, we additionally report the negative log-likelihood values versus the wall-clock time. The left panel shows the results when the parameters estimated by exact GPs are used, while the right panel presents the results when these parameters are doubled. Except for tapering, all approximations achive a similar log-likelihood for large runtimes. Vecchia approximations followed by full-scale approximations are the most accurate for a given computational budget.
Additional comparisons to exact calculations are presented in the Appendix H, where we report the RMSE between exact and approximate predictive means and variances.
@@ 3.3 House price dataset
3.4 Laegern dataset
In the following, we report the results for the Laegern dataset. The tuning parameters used are listed in Table 4. Figure 15 show the RMSE, the log-score, and the CRPS computed on the interpolation and extrapolation test sets. We again find that Vecchia approximations are the most accurate ones even when few neighbors are used. No other approximation reaches the accuracy of Vecchia approximations on the interpolation test set not even for long runtimes. For the extrapolation test data set, the results of the Vecchia approximation, the full-scale approximation, SPDE and the MRA are very close for large runtimes. Overall, FRK and FITC approximations are the least accurate approximations with poor scores even for large runtimes. Note that some very bad results for FITC and FRK are not shown in Figure 15 for better visibility of the other results, but they can be found in Appendix I. Unfortunately, it was not possible to consider results from periodic embedding as its run time exceeds our 3-hours limit. In addition, Figure 16 shows the negative log-likelihood evaluated at the parameter estimates from a Vecchia approximation with the largest considered number of neighbors on the left and two times these values on the right. Vecchia approximations achieves the lowest negative log-likelihood with both sets of covariance parameters.
Figure 15: RMSE, log-score, and CRPS on the Laegern dataset.
@@ 3.4 Laegern dataset
3.5 MODIS 2016 dataset
Next, we report the results for the MODIS 2016 dataset. The choices of the tuning parameters for all approximations are displayed in Table 5. Note that for the SPDE approach, a very limited range of approximation complexity can considered due to the algorithm’s failure to converge when the largest triangulation edge is smaller than $0.36$. Figures 17 presents the prediction accuracy results versus the wall-clock-time including the RMSE, the log-score, and the CRPS. Once again, the Vecchia approximation emerges as the most accurate approximation, excelling already for short runtimes, followed by the MRA. Periodic embedding also achieves good results in terms of RMSE and CRPS, but shows a relatively higher log-score due to very small estimated variances. In Figure 18, we plot the negative log-likelihood evaluated at the parameter estimates from a Vecchia approximation with the largest considered number of neighbors on the left and two times these values on the right. Vecchia approximations quickly achieve much lower negative log-likelihoods compared to all other methods. Note that negative log-likelihood values computed with the MRA have not been included in Figure 18 as they are very large, but they can be found in Tables 79 and 80 in Appendix J.
Figure 17: Test RMSE, log-score, and CRPS on the MODIS 2016 dataset.
@@ 3.5 MODIS 2016 dataset
3.6 MODIS 2023 dataset
We next discuss the results for the MODIS 2023 dataset. The choices of the tuning parameters for all approximations are shown in Table 6. Due to the large data size, for the MRA it was only possible to implement one level of complexity, corresponding to $r=1$. In addition, for FRK we had to leave out the case with 3 resolutions as this exceeds our time limits as well. Figure 19 reports the test set prediction accuracy results. Similarly as for other datasets, the Vecchia approximation quickly converges do a certain value. Interestingly, the full-scale and FITC approximations attain better prediction accuracy measures than Vecchia approximations. Figure 20 displays the negative log-likelihoods versus time. Vecchia approximations converge quickly to low negative log-likelihoods while all other methods fail to converge. Note that the results of MRA are absent from the plot as they are very large, but they can be found in Tables 82 and 83.
Figure 19: Test RMSE, log-score, and CRPS on the MODIS 2023 dataset.
@@ 3.6 MODIS 2023 dataset
3.7 Results without multi-core parallelization
To analyze the impact of multi-core parallelization, we repeat the analysis for the large simulated datasets with $N=100^{\prime}000$ and the house price dataset using a machine with single-core processor and the same processor model, clock speed, and amount of RAM. The results for the simulated datasets with ranges $\rho=0.05$, $\rho=0.2$, and $\rho=0.5$ can be found in Appendices L, M, and N, respectively. In Appendix O, we report the results for the house price data sets. The results are overall very similar to the results when using multi-core parallelization. Vecchia approximations are the most accurate approximations for a given computational budget in almost all settings. We note that there differences among the methods regarding the speed-up of multi-core parallelization. However, given that approximations such as Vecchia, FITC, full-scale, and tapering are implemented in a different library compared to MRA and SPDE approximations, it is difficult to say whether these speed-up differences are method-inherent or particularities of the specific implementations.
4 Conclusion
We conduct an extensive accuracy-runtime trade-off comparison between eight popular Gaussian process approximations on simulated data with different sample sizes and large-scale real-world spatial datasets. We analyze the accuracy of the different methods concerning likelihood evaluation, parameter estimation, and prediction taking into account the time required to achieve a certain accuracy. Overall, Vecchia approximations clearly emerge as the most accurate approximation type. There are several open questions for future research. First, we only considered the regression case. Future research can analyze the quality of Gaussian process approximations using non-Gaussian likelihoods. For this, a challenge is that there are no closed-form expressions for marginal likelihoods and predictive distributions which means that additional approximations have to be made. Other covariance functions besides Matérn covariance functions can also be considered. Furthermore, while we did consider eight of the most prominent Gaussian process approximations in spatial statistics, we obviously did not include all available approximations. Future research can extend this to more methods.
Acknowledgments
Appendix A Supplementary material
Appendix B Simulated data with $N=10^{\prime}000$ and practical range of $0.2$
B.1 Log-likelihood evaluation

### v2

@@ 2.4 GP approximations and choice of tuning parameters
We consider the following approximation techniques: Vecchia approximations, covariance tapering, modified predictive process aka Fully Independent Training Conditional (FITC) approximations, full-scale approximations (FSA) combining predictive processes with tapering, multi-resolution approximations (MRA), the stochastic partial differential equations (SPDE)-based approximation of Lindgren et al. (2011), fixed rank kriging (FRK), and the periodic embedding approach of Guinness (2019).
@@
The runtime and accuracy of the different approximation methods are determined by one or several tuning parameters such as the number of neighbors for Vecchia approximations and the number of knots for modified predictive process approximations. For every approximation, we use up to five different choices of tuning parameters. The selection of tuning parameter is done such that the different methods have roughly similar run-times. The specific choices of the tuning parameters thus vary across data sets and are reported in Appendix A.2. In the simulated experiments, the choice of the tuning parameters is done based on the time for evaluating the log-likelihood. In real-world applications, we instead use the total time for estimation and prediction on the (“interpolation”) test set. We also set an upper bound on the total runtime for every iteration in the experiments, including likelihood evaluation, estimation, and prediction, and we do not consider tuning parameter choices that exceed this limit. For simulated data sets with sample sizes $N=10^{\prime}000$ and $N=100^{\prime}000$, this limit is $10$ and $30$ minutes per repetition, respectively. For the real-world data, we use the following limits: $30$ minutes for the house price data set, two hours for the MODIS 2016 data set, three hours for the Laegern data set, and four hours for the MODIS 2023 data set.
@@
Vecchia approximations are applied to the observable process $y(\cdot)$ and not the latent process $f(\cdot)$, random orderings are used as this has been found to give accurate approximations (Guinness, 2018), nearest neighbors are determined using the Euclidean distance, and for prediction, the observed locations appear first in the ordering and we condition on the observed locations only. For FITC and FSA approximations, the knots are determined using the kmeans++ algorithm. For covariance tapering, the time for log-likelihood evaluation does not include the time required to perform a symbolic Cholesky decomposition for sparsity pattern since this needs to be done only once.
@@ 2.5 Software and hardware resources
We use the following software implementations for conducting our experiments. For Vecchia approximations, covariance tapering, FITC approximations, and full-scale approximations, we use the GPBoost library version 1.5.1 (Sigrist et al., 2021). All these methods are implemented in C++ in a comparable manner which makes their comparison as fair as possible. For the other four techniques, we use the following R packages: GPvecchia version 0.1.5 for the multi-resolution approximation, INLA version 24.6.27 and rSPDE version 2.3.3 for the SPDE approach, FRK version 2.3.0 for fixed rank kriging, and npspec version 0.1.0 for periodic embedding (https://github.com/joeguinness/npspec). Parameter estimation is done with the default optimizers in the corresponding software packages. This means that differences in the estimation times of the methods implemented in GPBoost and of the remaining methods are also influenced by the optimizers used. This is, arguably, not ideal, but it is not possible to use the same optimizers in all packages. However, the runtimes for log-likelihood evaluation for the simulated and real-world data sets and for prediction for the simulated data sets do not suffer from this drawback.
@@ 3.1 Simulated data with $N=10^{\prime}000$
We first discuss the results for the simulated data with a sample of size $N=10^{\prime}000$ and a practical range of $0.2$. We use figures to present the results in the main article, the results in tabular form can be found in the supplementary material. The tuning parameters used for every method are reported in Table 1 in Appendix A.2.
@@
In Figure 1, we show the absolute mean difference between the exact and approximate log-likelihood versus the evaluation time. The left panel shows the results when the evaluation is done at the true data-generating parameters, while the right panel shows the results using “wrong” parameters obtained by doubling the true parameters. Vecchia approximations clearly achieve the highest accuracy. The difference to the exact log-likelihood decreases quickly with respect to time, approaching values close to zero even for small numbers of neighbors. The second most accurate approximation for a given computational budget is the FITC approximation followed by the MRA and FSA. Tapering is the method with the overall lowest accuracy, achieving a high accuracy only for relatively large taper ranges. The results for larger and smaller practical ranges of $0.5$ and $0.05$ are reported in Appendix B.2 and Appendix B.3, respectively. Overall, we observe similar findings. Vecchia approximations are clearly the most accurate approximations. As expected, FITC approximations are more accurate for large range parameters compared to small ones, and vice versa for covariance tapering. Note that the SPDE-based approach is not included in the log-likelihood comparison as, to the best of our knowledge, the corresponding software implementation does not allow for evaluating the log-likelihood at user-defined covariance parameters. Further, FRK and and periodic embedding do not try to approximate a GP with a Matern covariance function, and it is thus pointless to analyze the accuracy of log-likelihood evaluations.
@@
Figure 1: Absolute mean difference between approximate and exact log-likelihood on simulated data sets with a practical range of $0.2$ of and a sample size $N=10^{\prime}000$. The true data-generating parameters are used on the left and “wrong” parameters obtained by doubling the true ones are used on the right.
@@
We next present the results for parameter estimation. In Figure 2, we show the bias and the MSE for the range, marginal variance, and error variance parameters, respectively, versus the time required to estimate these parameters. We again find that Vecchia approximations are clearly the most accurate with the lowest bias and MSE for most computational times and very small biases and MSEs even for small estimation times. The range and the variance parameters are overestimated by most methods. For tapering and FITC, some results with extreme values are excluded from the plots for better visibility of the other results. The full results in tabular form can be found in Appendix B.1. Note that we were not able to extract estimates for range and marginal variance parameters using the corresponding software implementation for the SPDE approach.
@@
Figure 2: Bias and MSE of the estimates for the range, marginal variance, and error term variance on simulated data sets with a practical range of $0.2$ and a sample size $N=10^{\prime}000$. The dashed lines correspond to the results from the exact calculations.
@@
In Figure 3, we report the prediction accuracy results including RMSEs, log-scores, and KL divergences to exact calculations for predicting the latent process at training locations, the “interpolation” test data, and the “extrapolation” test locations. We observe that Vecchia approximations are clearly the most accurate compared to all other methods regarding all evaluation metrics for most computational times. Overall, the MRA is the second most accurate method in this experiment, followed by the FITC and SPDE approximations. The FITC, FSA, MRA, and SPDE methods eventually achieve similar prediction accuracies as Vecchia approximations but for much larger computational times. We were not able to attain this level of accuracy with tapering in the computational times considered. Finally, FRK results in considerably less accurate predictions. For the “extrapolation” test set, we observe that the variability is much higher for all methods due to the uncertainty involved in making predictions outside of the training domain. Note that for FRK, the metrics calculated on the “extrapolation” set are not included in the figure as they were very large, i.e., very inaccurate. All results in tabular form can be found in Appendix B.1. In Appendix B.1, we additionally report the RMSE comparing the approximate predictive means and variance to the exact values.
@@
Figure 3: Average RMSE, log-score, and KL divergence between exact and approximate predictions on simulated data sets with a practical range of $0.2$ and a sample size $N=10^{\prime}000$. We distinguish between predicting the latent process at (i) training locations (left), (ii) interpolation test locations (center), and (iii) extrapolation locations (right). The dashed lines correspond to the results from the exact calculations.
The results for experiments conducted on simulated data sets with practical ranges of 0.5 and 0.05 are reported in Appendices B.2 and B.3, respectively. In both scenarios, Vecchia approximations consistently rank among the most accurate, particularly for short runtimes. The MRA shows similar behavior across the different ranges, though it tends to be slightly less accurate than Vecchia approximations. As expected, all techniques generally are more accurate for a larger range parameter. An exception is covariance tapering, which, also as expected, is more accurate for small range parameters and less accurate for large ones.
3.2 Simulated data with $N=100^{\prime}000$
Next, we discuss the results for the simulated data with a sample of size $N=100^{\prime}000$. The tuning parameters used for every method are reported in Table 2 in Appendix A.2. Figure 16 in Appendix A.1 displays the average approximate negative log-likelihood versus the evaluation time for a practical range of $0.2$. Recall that, for this large sample size, we do not consider the difference to the exact log-likelihood, as the latter can not be calculated in a reasonable amount of time. The results are very similar to the ones of the smaller sample size. The Vecchia approximation is again the most accurate method, with the maximal accuracy being achieved for relatively small runtimes.
Concerning parameter estimation accuracy, Figure 17 in Appendix A.1 reports the bias and MSE for the range, marginal variance, and error variance parameters, respectively, versus the time required to estimate these parameters. We again find very similar results as for the smaller sample size of $N=10^{\prime}000$. Vecchia approximations are clearly the most accurate with the lowest bias and MSE for most computational times and very small biases and MSEs even for small number of neighbors. Note that for tapering, some results with extreme values are again excluded from the plots for better visibility of the other results. The full results in tabular form can be found in Appendix B.4.
For assessing prediction accuracy, Figure 18 in Appendix A.1 shows the RMSE and log-scores for predicting the latent process at training locations, the “interpolation” test data, and the “extrapolation” test data. We find that Vecchia approximations are the most accurate compared to all other methods regarding all evaluation metrics for most computational times. The SPDE-based approach results in the second most accurate predictions overall followed by the MRA and the FITC approximation. FRK and tapering generally generate the least accurate predictions. Note that the FITC approximation performs relatively well on the extrapolation test data. For FRK, the results have been excluded again from the extrapolation plots due to extreme and very inaccurate values; they can be found in Appendix 49 in tabular form. The results for the experiments using simulated data with $N=100^{\prime}000$ and practical ranges of $0.5$ and $0.05$ are reported in Appendix B.5 and Appendix B.6, respectively. Overall, the results are qualitatively very similar to the above-reported results.
3.3 House price data set
In this subsection, we present the results for the house price data set. Note that we do not use the periodic embedding approach for this data set, as the locations are not arranged on a lattice. The tuning parameters used on this data set are reported in Table 3 in Appendix A.2.
In Figure 4, we show the prediction accuracy results on the “interpolation” and “extrapolation” test sets of the housing data set. Specifically, we plot the RMSE, the log-score and the CRPS versus the prediction time. In addition, we also report the results when doing predictions with exact calculations. To further compare the differences between exact and approximate calculations, Figure 5 shows the KL divergence between the exact and approximate predictive distributions versus the runtime on the two test sets. We observe that Vecchia approximations are again the most accurate approximations having almost equal prediction accuracy as exact calculations on both the interpolation and the extrapolation test sets even for small runtimes. This is confirmed by the very small KL divergences in Figure 5 of predictive distributions of Vecchia approximations compared to exact calculations. For the interpolation test set, full-scale approximations are the second most accurate. For long runtimes, the prediction accuracy measures are similar across all approximations except for FRK. Note that for the extrapolation test data, the FITC and full-scale approximations with a small number of inducing points lead to more accurate predictions than an exact Gaussian process models. As above, the poor FRK scores on the “extrapolation” set had to be removed from the plots, but can be found in Appendix B.7. In Figure 19 in Appendix A.1, we additionally report the negative log-likelihood values versus the wall-clock time. Except for tapering, all approximations yield a similar log-likelihood for large runtimes. Vecchia approximations followed by full-scale approximations are the most accurate for a given computational budget. Additional comparisons to exact calculations are presented in Appendix B.7, where we report the RMSE between exact and approximate predictive means and variances.
3.4 Laegern data set
In the following, we report the results for the Laegern data set. The tuning parameters used are listed in Table 4 in Appendix A.2. Figure 6 show the RMSE, the log-score, and the CRPS computed on the interpolation and extrapolation test sets. We again find that Vecchia approximations are the most accurate ones even when few neighbors are used. No other approximation reaches the accuracy of Vecchia approximations on the interpolation test set not even for long runtimes. For the extrapolation test data set, the results of the Vecchia approximation, the full-scale approximation, SPDE and the MRA are very close for large runtimes. Overall, FRK and FITC approximations are the least accurate approximations with poor scores even for large runtimes. Note that some very bad results for FITC and FRK are not shown in Figure 6 for better visibility of the other results, but they can be found in Appendix B.8. Unfortunately, it was not possible to consider results from periodic embedding as its run time exceeds our 3-hours limit. In addition, Figure 20 in Appendix A.1 shows the negative log-likelihood versus the runtime. Vecchia approximations clearly converges the fastest and achieves the lowest negative log-likelihood.
Figure 6: RMSE, log-score, and CRPS on the Laegern data set.
3.5 MODIS 2016 data set
Next, we report the results for the MODIS 2016 data set. The choices of the tuning parameters for all approximations are reporte in Table 5 in Appendix A.2. Note that for the SPDE approach, a limited range of approximation complexity can considered due to the algorithm’s failure to converge when the largest triangulation edge is smaller than $0.36$. Figures 7 presents the prediction accuracy results versus the wall-clock-time including the RMSE, the log-score, and the CRPS. Once again, the Vecchia approximation emerges as the most accurate approximation, excelling already for short runtimes, followed by the MRA. Periodic embedding also achieves good results in terms of RMSE and CRPS, but shows a relatively higher log-score due to very small estimated variances. In Figure 8, we plot the negative log-likelihood versus the wall-clock time. Vecchia approximations quickly achieve much lower negative log-likelihoods compared to all other methods. Note that negative log-likelihood values computed with the MRA have not been included in Figure 8 as they are very large, but they can be found in Tables 79 and 80 in Appendix B.9.
Figure 7: Test RMSE, log-score, and CRPS on the MODIS 2016 data set.
3.6 MODIS 2023 data set
We next discuss the results for the MODIS 2023 data set. The choices of the tuning parameters for all approximations are reported in Table 6 in Appendix A.2. Due to the large data size, it was only possible to implement one level of complexity, $r=1$, for the MRA. In addition, for FRK we had to leave out the case with 3 resolutions as this exceeds our time limits as well. Figure 9 reports the test set prediction accuracy results. Similarly as for other data sets, the Vecchia approximation quickly converges do a certain value. The full-scale and FITC approximations attain better prediction accuracy measures compared Vecchia approximations. This is likely a consequence of the fact that the test data contains a lot of extrapolation areas. It was already observed in Section 3.3 that low-rank approximations can be more accurate in extrapolation than exact calculations. Figure 10 displays the negative log-likelihoods versus time. Vecchia approximations converge quickly to low negative log-likelihoods while all other methods fail to converge. Note that the results of MRA are absent from the plot as they are very large, but they can be found in Tables 82 and 83.
Figure 9: Test RMSE, log-score, and CRPS on the MODIS 2023 data set.
3.7 Results without multi-core parallelization
To analyze the impact of multi-core parallelization, we repeat the analysis for the large simulated data sets with $N=100^{\prime}000$ and the house price data set using a machine with single-core processor and the same processor model, clock speed, and amount of RAM. The results for the simulated data sets with ranges $\rho=0.05$, $\rho=0.2$, and $\rho=0.5$ can be found in Appendices B.11, B.12, and B.13, respectively. In Appendix B.14, we report the results for the house price data sets. The results are overall very similar to the results when using multi-core parallelization. Vecchia approximations are the most accurate approximations for a given computational budget in almost all settings. We note that there are differences among the methods regarding the speed-up of multi-core parallelization. However, given that approximations such as Vecchia, FITC, full-scale, and tapering are implemented in a different library compared to MRA and SPDE approximations, it is difficult to say whether these speed-up differences are method-inherent or due the specific implementations.
4 Conclusion
We conduct an extensive accuracy-runtime trade-off comparison between eight popular Gaussian process approximations on simulated data with different sample sizes and multiple large-scale real-world spatial data sets. We analyze the accuracy of the different methods concerning likelihood evaluation, parameter estimation, and prediction taking into account the time required to achieve a certain accuracy. Overall, Vecchia approximations clearly emerge as the most accurate approximation type. There are several open questions for future research. First, we only considered the regression case. Future research can analyze the quality of Gaussian process approximations using non-Gaussian likelihoods. For this, a challenge is that there are no closed-form expressions for marginal likelihoods and predictive distributions which means that additional approximations have to be made. Other covariance functions besides Matérn covariance functions can also be considered. Furthermore, while we did consider eight of the most prominent Gaussian process approximations in spatial statistics, we obviously did not include all available approximations. Future research can extend this to more methods.
Acknowledgments
Appendix
A.1 Additional figures
[…]
Figure 17: Bias and MSE of the estimates for the range, marginal variance, end error term variance on simulated datasets with a practical range of $0.2$ and a sample size N=100’000,
Figure 18: Average RMSE and log-score on simulated datasets with a practical range of $0.2$ and a sample size N=100’000.
A.2 Tuning parameters
@@
Supplementary material
B.1 Simulated data with $N=10^{\prime}000$ and practical range of $0.2$
B.1.1 Log-likelihood evaluation